# On the cross-correlation properties of large-size families of Costas arrays

### v1

@@ III The family of Costas arrays $\mathcal{W}^{el}_{p}$ and its cross-correlation
The main goal of this section is to construct families of Costas arrays of the same order, whose sizes are as large as possible, and the cross-correlation values between any pair of Costas arrays in the families for any shifts are bounded by a value much smaller than the order [7]. Since the logarithmic Welch Costas arrays by Construction 2 are distinct from those by Construction 1 for $p>5$ [19], we extend the family of exponential Welch Costas arrays $\mathcal{W}_{p}$ by including the family of logarithmic Welch Costas arrays and study the cross-correlation properties of this larger family. Note that the size of this family is twice as large as the family $\mathcal{W}_{p}$.
@@
For $p>5$, the family $\mathcal{W}^{el}_{p}$ of Costas arrays of order $(p-1)$ is defined as
@@
As shown in Table III, we list exhaustively the maximal cross-correlation values of two Costas arrays in the family $\mathcal{W}_{p}^{el}$, i.e., $\mathcal{C}(\mathcal{W}^{el}_{p})=\max\limits_{(u,v)}\max\limits_{f\neq g}\Psi_{f,g}(u,v)$, where $f,g\in\mathcal{W}_{p}^{el}$ and $20\leq|u|,|v|\leq p-$ 2, for $7\leq p\leq 277$. The maximal cross-correlation values of the family $\mathcal{W}_{p}$ are also included in Table III for comparison, denoted by $\mathcal{C}(\mathcal{W}_{p})$, which can also be found in [7].
@@
Recall that the maximal cross-correlation of $\mathcal{W}_{p}$, for arbitrary $u$ and $v=0$, is equal to $\frac{p-1}{t}$, where $t$ is the smallest prime divisor of $\frac{p-1}{2}$ and, for arbitrary $u$ and $v\neq 0$, is bounded by $1+\left\lfloor\left(1-\frac{2}{p-1}\right)p^{1/2}\right\rfloor$. It then follows that when $p$ is not a safe prime, $\mathcal{C}(\mathcal{W}_{p})=\frac{p-1}{t}$, since $1+\left\lfloor\left(1-\frac{2}{p-1}\right)p^{1/2}\right\rfloor\leq\frac{p-1}{t}$ and when $p$ is a safe prime, $\mathcal{C}(\mathcal{W}_{p})\leq 1+\left\lfloor\left(1-\frac{2}{p-1}\right)p^{1/2}\right\rfloor$, since $\frac{p-1}{t}=2$.
It is observed that there are also two cases on the maximal cross-correlation of the family $\mathcal{W}_{p}^{el}$: (1) When $p$ is not a safe prime, $\mathcal{C}(\mathcal{W}^{el}_{p})=\mathcal{C}(\mathcal{W}_{p})=\frac{p-1}{t}$, where $t$ is the least prime divisor of $\frac{p-1}{2}$. For example, let $p=103$ and factorize $p-1$ to obtain $103-1=2\times 3\times 17$, writing the factors in the ascending order. The least prime divisor of $\frac{p-1}{2}$ is equal to $3$, then $\frac{p-1}{t}=\frac{102}{3}=34$, corresponding to the result in the column $\mathcal{C}(\mathcal{W}^{el}_{p})$; (2) When $p$ is a safe prime, $\mathcal{C}(\mathcal{W}^{el}_{p})\neq\mathcal{C}(\mathcal{W}_{p})$.
To prove some of the above observations, we need to calculate the cross-correlation between any two Costas arrays $f,g$ in $\mathcal{W}^{el}_{p}$ at the horizontal shift $u$ and the vertical shift $v$. There are three cases where $f$ and $g$ are both exponential Welch Costas arrays, both logarithmic Welch Costas arrays or one is an exponential Welch Costas array while the other is a logarithmic Welch Costas array. Since the maximal cross-correlation between either two exponential Welch Costas arrays or two logarithmic Welch Costas arrays already have settled bounds [8], the remaining question is to compute the cross-correlation between an exponential Costas array and a logarithmic one at the shift $(u,v)$.
@@
It seems hard to directly determine the number of solutions of Eq. (2) because this is a transcendental equation system over a finite field. Next, we give an upper bound of the number of solutions of Eq. (2) for arbitrary $u$ and $v$, using the idea from [11] to simplify the equation system into a polynomial equation.
@@
Suppose that $(x_{1},y_{1}),(x_{2},y_{2}),\cdots,(x_{N},y_{N})$ are the solutions of Eq. (2) where $10\leq x_{i},y_{i}\leq p-$ 1 for $i=1,2,\cdots,N$. Then, for $i=1,2,\cdots,N$, we have
@@
Noting that $y_{i}-v,y_{i+1}-v\in\{0,1,\ldots,p-1\}$ and using  [11, Lemma 1], we have
@@
Taking $z$ as $\frac{1}{2}\log_{\alpha}p$ which is a logarithmic function on real numbers, we have
@@
$\mathcal{C}(\mathcal{W}^{el}_{p})\leq\begin{cases}\max\left\{4p\,\log_{p}\alpha+1,1+\left\lfloor\left(1-\frac{2}{p-1}\right)p^{1/2}\right\rfloor\right\},&\text{ if }p\text{ is a safe prime, }\\ \max\left\{4p\,\log_{p}\alpha+1,\frac{p-1}{t}\right\},&\text{ otherwise, }\end{cases}$ ($\ast$)
@@
This result directly follows from Lemma 4 and [8, Theorem 1]. ∎
@@
$p$ $\alpha$ Bound ()(*)( ∗ ) $p$ $\alpha$ Bound ()(*)( ∗ ) $p$ $\alpha$ Bound ()(*)( ∗ )
@@ IV Bounds on the maximal cross-correlation of $\mathcal{P}_{p}$ and $\mathcal{PW}_{p}$
The extended family $\mathcal{PW}_{p}$ investigated by Ardalani [10], which incorporates the family of exponential Welch Costas arrays and the family of power permutations, is defined as follows.
@@
[10] For prime $p\geq 5$, the family $\mathcal{PW}_{p}$ of order $p-1$, is defined as
@@
where $\mathcal{W}_{p}$ and $\mathcal{P}_{p}$ are defined in Construction 3 and Construction 5, respectively. The size of the family $\mathcal{PW}_{p}$ equals $12\phi(p-1)-$ 1, where $\phi$ denotes Euler’s totient function.
@@
Suppose that there exists an integer $d$ with $\operatorname{gcd}(d,p-1)=1$, and a primitive element $\alpha$ satisfting $x^{d}\equiv\alpha^{x-1}\bmod p$ for all $x\in\mathbb{F}^{*}_{p}$. Then we have
@@
Via some substitutions, we have
@@
Ardalani [10] used exhaustive search to determine the maximal cross-correlation values of the family $\mathcal{P}_{p}$. The maximal cross-correlation value is defined as $\mathcal{C}\left(\mathcal{P}_{p}\right)=\max\limits_{(u,v)}\max\limits_{f,g\in\mathcal{P}_{p}}\Psi_{f,g}(u,v)$ where $(u,v)\neq(0,0)$ if $f=g$ and $20\leq|u|,|v|\leq p-$ 2 otherwise. Similarly, Ardalani [10] computed the maximal cross-correlation values for two elements from the family $\mathcal{PW}_{p}$, i.e., $\mathcal{C}\left(\mathcal{PW}_{p}\right)=\max\limits_{(u,v)}\max\limits_{f,g\in\mathcal{PW}_{p}}\Psi_{f,g}(u,v)$, where $(u,v)\neq(0,0)$ if $f=g$ and $20\leq|u|,|v|\leq p-$ 2 otherwise. These computations were performed for all primes $5\leqslant p\leqslant 277$. The maximal cross-correlation values of the family $\mathcal{W}_{p}$ are included as well for comparison, see Table V.
@@
TABLE V: Maximal Cross-correlation of $\mathcal{W}_{p}$, $\mathcal{P}_{p}$ and $\mathcal{PW}_{p}$ [10]
For the maximal cross-correlation of the family of powering permutations $\mathcal{P}_{p}$, Ardalani [10] settled the two cases when $u=0,v=0$ and $u\neq 0,v=0$, by some counting techniques and character theory. We now deal with the case where $u=0,v\neq 0$ via Kelley and Owen’s result [18]. However, it seems still difficult to solve the general case when $u\neq 0,v\neq 0$, which is left as an open problem.
Let $f(x)=x^{d_{1}}$ and $g(x)=x^{d_{2}}$ be two power permutations over $\mathbb{F}_{p}$, where $x\in\mathbb{F}_{p}\backslash\{0\}$, $\operatorname{gcd}(d_{1},p-1)=1$ and $\operatorname{gcd}(d_{2},p-1)=1$. Then, the formulation of the cross-correlation of $f,g$ at the shift $(u,v)$ is the following:
$\Psi_{f,g}(u,v)=|\{x\in\mathbb{F}_{p}\backslash\{0\}|x^{d_{1}}\bmod p+v=(x+u)^{d_{2}}\bmod p\}|.$ (6)
@@
$\max_{v\neq 0}\Psi_{f,g}(0,v)\leq\frac{1}{2}+\sqrt{p-1}.$
@@
With respect to the family $\mathcal{PW}_{p}$, we give the first nontrivial upper bound of the maximal cross-correlation for arbitrary $u$ and $v=0$. Given that there are some upper bounds of the maximal cross-correlation of $\mathcal{W}_{p}$ and $\mathcal{P}_{p}$, it is crucial to get upper bound between an exponential Welch Costas array and a power permutation. Let $f(x)=x^{d}$, where $x\in\mathbb{F}_{p}\backslash\{0\}$ and $\operatorname{gcd}(d,p-1)=1$, be a power permutation over $\mathbb{F}_{p}$, and let $g=W_{1}^{\exp}(p,\alpha,0)$ be an exponential Welch Costas array. Then, the formulation of the cross-correlation between a power permutation and an exponential Welch Costas array at $(u,v)$ is [10]:
$\Psi_{f,g}(u,v)=|\{x\in\mathbb{F}_{p}\backslash\{0\}|x^{d}\bmod p+v=\alpha^{x-1+u}\bmod p\}|.$ (9)
@@
$\Psi_{f,g}(u,0)=|\{x\in\mathbb{F}_{p}\backslash\{0\}|x^{d}\bmod p=\alpha^{x-1+u}\bmod p\}|.$
To proceed, let $e$ denote the modular inverse of $d$ such that $de\equiv 1(\bmod\enspace p-1)$. By raising both sides of the congruence $x^{d}\bmod p=\alpha^{x-1+u}\bmod p$ to the power of $e$, we obtain
@@
Note that $\alpha^{e}$ is still a primtive element over $\mathbb{F}_{p}$ and define $\beta=\alpha^{e}$. Substituting $\beta$ into the equation, we rewrite it as
@@
Observe that the function $f=\beta^{x-1+u},\text{ where }x\in[p-1]$ generates an exponential Costas array $W_{1}^{\exp}(p,\beta,u)$. The equation $x=\beta^{x-1+u}\bmod p$ implies that $x$ is a fixed point on the main diagonal of this Costas array. Consequently, the number of solutions is equal to the number of fixed points on the main diagonal. Since the fixed points on the main diagonal form a Golomb ruler, via Lemma 1, the number of the fixed points is strictly less than
@@
Case (i): Suppose that $f$ and $g$ are both exponential Costas arrays. According to [7, Theorem 1], we have
$\max\limits_{u}\max\limits_{\begin{subarray}{c}f,g\in\mathcal{W}_{p}\subset\mathcal{PW}_{p}\end{subarray}}\Psi_{f,g}(u,0)=\frac{p-1}{t}.$
Case (ii): Suppose that $f$ and $g$ are both power permutations. From [10, Theorem 5.7 and Theorem 5.9], it follows that, for $(u,v)=(0,0)$,
$\max\limits_{u}\max\limits_{\begin{subarray}{c}f\neq g\end{subarray}}\Psi_{f,g}(0,0)=\frac{p-1}{t},$
@@
$\max\limits_{u}\max\limits_{\begin{subarray}{c}f\neq g\end{subarray}}\Psi_{f,g}(0,0)\leq\left\lceil\frac{p-2}{p-1}(1+\sqrt{p})\right\rceil.$
@@
$\max\limits_{u}\max\limits_{\begin{subarray}{c}f,g\in\mathcal{W}_{p}\subset\mathcal{PW}_{p}\end{subarray}}\Psi_{f,g}(u,0)=\frac{p-1}{t},$
since
$\frac{p-1}{t}\geq\sqrt{2(p-1)}\geq\frac{p-2}{p-1}(1+\sqrt{p}).$
@@
and the numerical results in Table V.
@@
For a prime $p\geq 5$, the family $\mathcal{PW}^{l}_{p}$ of order $p-1$, is defined as
@@
where $\mathcal{W}^{l}_{p}$ and $\mathcal{P}_{p}$ are defined in Construction 4 and Construction 5, respectively. The size of the family $\mathcal{PW}^{l}_{p}$ equals $12\phi(p-1)-$ 1, where $\phi$ is Euler’s totient function.

### v2

@@ III The family of Costas arrays $\mathcal{W}^{el}_{p}$ and its cross-correlation
The main goal of this section is to extend the family of exponential Welch Costas arrays $\mathcal{W}_{p}$ by including the family of logarithmic Welch Costas arrays and study the cross-correlation properties of this larger family since the logarithmic Welch Costas arrays by Construction 2 are distinct from those by Construction 1 for $p>5$ [DG2009Symmetry]. To best of our knowledge, this new family has not been investigated before. Note that the size of this family is twice as large as the family $\mathcal{W}_{p}$ which has been widely studied.
@@
For $p>5$, the family $\mathcal{W}^{el}_{p}$ of Costas arrays of order (p1)(p-1) is defined as
@@
As shown in Table III, we list exhaustively the maximal cross-correlation values of two Costas arrays in the family $\mathcal{W}_{p}^{el}$, i.e., $\mathcal{C}(\mathcal{W}^{el}_{p})=\max\limits_{(u,v)}\max\limits_{f\neq g}\Psi_{f,g}(u,v)$, where $f,g\in\mathcal{W}_{p}^{el}$ and $0\leq|u|,|v|\leq p-2$, for $7\leq p\leq 277$. The maximal cross-correlation values of the family $\mathcal{W}_{p}$ are also included in Table III for comparison, denoted by $\mathcal{C}(\mathcal{W}_{p})$, which can also be found in [DG2011Maximal].
@@
Recall that the maximal cross-correlation of $\mathcal{W}_{p}$, for arbitrary $u$ and $v=0$, is equal to $\frac{p-1}{t}$, where $t$ is the smallest prime divisor of $\frac{p-1}{2}$ and, for arbitrary $u$ and $v\neq 0$, is bounded by $1+\left\lfloor\left(1-\frac{2}{p-1}\right)p^{1/2}\right\rfloor$. It then follows that when $p$ is not a safe prime, $\mathcal{C}(\mathcal{W}_{p})=\frac{p-1}{t}$, since $1+\left\lfloor\left(1-\frac{2}{p-1}\right)p^{1/2}\right\rfloor\leq\frac{p-1}{t}$ and when $p$ is a safe prime, $\mathcal{C}(\mathcal{W}_{p})\leq 1+\left\lfloor\left(1-\frac{2}{p-1}\right)p^{1/2}\right\rfloor$, since $\frac{p-1}{t}=2$. Here, a prime $p$ is a safe prime if and only if $p=2r+1$ and $r$ is also a prime (also known as Sophie Germain prime).
It is observed that there are also two cases on the maximal cross-correlation of the family $\mathcal{W}_{p}^{el}$: (1) When $p$ is not a safe prime, see $p$ values with no underline in Table III, $\mathcal{C}(\mathcal{W}^{el}_{p})=\mathcal{C}(\mathcal{W}_{p})=\frac{p-1}{t}$, where $t$ is the least prime divisor of $\frac{p-1}{2}$. For example, let $p=103$ and factorize p1p-1 to obtain $103-1=2\times 3\times 17$, writing the factors in the ascending order. The least prime divisor of $\frac{p-1}{2}$ is equal to $3$, then $\frac{p-1}{t}=\frac{102}{3}=34$, corresponding to the result in the column $\mathcal{C}(\mathcal{W}^{el}_{p})$; (2) When $p$ is a safe prime, see underline $p$ values in Table III, $\mathcal{C}(\mathcal{W}^{el}_{p})\neq\mathcal{C}(\mathcal{W}_{p})$ and $\mathcal{C}(\mathcal{W}^{el}_{p}),\mathcal{C}(\mathcal{W}_{p})$ are both small compared to the order of the arrays.
To prove some of the above observations, we need to calculate the cross-correlation between any two Costas arrays $f,g$ in $\mathcal{W}^{el}_{p}$ at the horizontal shift $u$ and the vertical shift $v$. There are three cases where $f$ and $g$ are both exponential Welch Costas arrays, both logarithmic Welch Costas arrays or one is an exponential Welch Costas array while the other is a logarithmic Welch Costas array. Since the maximal cross-correlation between either two exponential Welch Costas arrays or two logarithmic Welch Costas arrays already have settled bounds [GW2020Note], the remaining question is to compute the cross-correlation between an exponential Costas array and a logarithmic one at the shift $(u,v)$.
@@
It seems hard to directly determine the number of solutions of Eq. (2) because this is a transcendental equation system over a finite field. Next, we give an upper bound of the number of solutions of Eq. (2) for arbitrary $u$ and $v$, using the idea from [GS2010Cycles] that simplifies the equation system into a polynomial equations.
@@
Suppose that $(x_{1},y_{1}),(x_{2},y_{2}),\cdots,(x_{N},y_{N})$ are the solutions of Eq. (2) where $0\leq x_{i},y_{i}\leq p-1$ for $i=1,2,\cdots,N$. Then, for $i=1,2,\cdots,N$, we have
@@
Noting that $y_{i}-v,y_{i+1}-v\in\{0,1,\ldots,p-1\}$ and using  [GS2010Cycles, Lemma 1], we have
@@
Taking $z$ as $\frac{1}{2}\log_{\alpha}p$ which is a logarithmic function on real numbers, we derive the upper bound on the number of solutions of Eq. (2)
@@
[…]
@@
This result directly follows from Lemma 4 and [GW2020Note, Theorem 1]. ∎
@@
$p$ $\alpha$ Bound ()(*) $p$ $\alpha$ Bound ()(*) $p$ $\alpha$ Bound ()(*)
@@ IV Bounds on the maximal cross-correlation of $\mathcal{P}_{p}$ and $\mathcal{PW}_{p}$
The extended family $\mathcal{PW}_{p}$ investigated by Ardalani [Ardalani2023Contributions, p. 117], which incorporates the family of exponential Welch Costas arrays and the family of power permutations, is defined as follows.
@@
[Ardalani2023Contributions] For prime $p\geq 5$, the family $\mathcal{PW}_{p}$ of order p1p-1, is defined as
@@
where $\mathcal{W}_{p}$ and $\mathcal{P}_{p}$ are defined in Construction 3 and Construction 5, respectively. The size of the family $\mathcal{PW}_{p}$ equals $2\phi(p-1)-1$, where $\phi$ denotes Euler’s totient function.
@@
Suppose that there exists an integer $d$ with $\operatorname{gcd}(d,p-1)=1$, and a primitive element $\alpha$ satisfying $x^{d}\equiv\alpha^{x-1}\bmod p$ for all $x\in\mathbb{F}^{*}_{p}$. Then we have
@@
By some substitutions, we have
@@
Ardalani [Ardalani2023Contributions] used exhaustive search to determine the maximal cross-correlation values of the family $\mathcal{P}_{p}$. The maximal cross-correlation value is defined as $\mathcal{C}\left(\mathcal{P}_{p}\right)=\max\limits_{(u,v)}\max\limits_{f,g\in\mathcal{P}_{p}}\Psi_{f,g}(u,v)$ where $(u,v)\neq(0,0)$ if $f=g$ and $0\leq|u|,|v|\leq p-2$ otherwise. Similarly, Ardalani [Ardalani2023Contributions] computed the maximal cross-correlation values for two elements from the family $\mathcal{PW}_{p}$, i.e., $\mathcal{C}\left(\mathcal{PW}_{p}\right)=\max\limits_{(u,v)}\max\limits_{f,g\in\mathcal{PW}_{p}}\Psi_{f,g}(u,v)$, where $(u,v)\neq(0,0)$ if $f=g$ and $0\leq|u|,|v|\leq p-2$ otherwise. These computations were performed for all primes $5\leqslant p\leqslant 277$. The maximal cross-correlation values of the family $\mathcal{W}_{p}$ are included as well for comparison, see Table V.
@@
TABLE V: Maximal Cross-correlation of $\mathcal{W}_{p}$, $\mathcal{P}_{p}$ and $\mathcal{PW}_{p}$ [Ardalani2023Contributions]
For the maximal cross-correlation of the family of power permutations $\mathcal{P}_{p}$, Ardalani [Ardalani2023Contributions] settled the two cases when $u=0,v=0$ and $u\neq 0,v=0$, by some counting techniques and character theory. We now deal with the case where $u=0,v\neq 0$ via Kelley and Owen’s result [KO2017trinomials]. However, it seems still difficult to solve the general case when $u\neq 0,v\neq 0$, which is left as an open problem.
Let $f(x)=x^{d_{1}}$ and $g(x)=x^{d_{2}}$ be two power permutations over $\mathbb{F}_{p}$, where $x\in\mathbb{F}_{p}^{*}$. Then, the formulation of the cross-correlation of $f,g$ at the shift $(u,v)$ is the following:
$\Psi_{f,g}(u,v)=|\{x\in\mathbb{F}_{p}^{*}|x^{d_{1}}\bmod p+v=(x+u)^{d_{2}}\bmod p\}|.$ (6)
@@
$\max_{v\neq 0}\max\limits_{f,g\in\mathcal{P}_{p}}\Psi_{f,g}(0,v)\leq\frac{1}{2}+\sqrt{p-1}.$
@@
With respect to the family $\mathcal{PW}_{p}$, we give the first nontrivial upper bound of the maximal cross-correlation for arbitrary $u$ and $v=0$. Given that there are some upper bounds of the maximal cross-correlation of $\mathcal{W}_{p}$ and $\mathcal{P}_{p}$, it is crucial to get upper bound between an exponential Welch Costas array and a power permutation. Let $f(x)=x^{d}$, be a power permutation over $\mathbb{F}_{p}$, and let $g=W_{1}^{\exp}(p,\alpha,0)$ be an exponential Welch Costas array. Then, the formulation of the cross-correlation between a power permutation and an exponential Welch Costas array at $(u,v)$ is [Ardalani2023Contributions]:
$\Psi_{f,g}(u,v)=|\{x\in\mathbb{F}_{p}^{*}|x^{d}\bmod p+v=\alpha^{x-1+u}\bmod p\}|.$ (9)
@@
$\Psi_{f,g}(u,0)=|\{x\in\mathbb{F}_{p}^{*}|x^{d}\bmod p=\alpha^{x-1+u}\bmod p\}|.$
To proceed, let $e$ denote the modular inverse of $d$ such that $de\equiv 1(\bmod\kern 5.0ptp-1)$. By raising both sides of the congruence $x^{d}\bmod p=\alpha^{x-1+u}\bmod p$ to the power of $e$, we obtain
@@
Note that $\alpha^{e}$ is still a primitive element over $\mathbb{F}_{p}$ and define $\beta=\alpha^{e}$. Substituting $\beta$ into the equation, we rewrite it as
@@
Observe that the function $f(x)=\beta^{x-1+u}\text{ where }x\in[p-1]$ generates an exponential Costas array $W_{1}^{\exp}(p,\beta,u)$. The equation $x=\beta^{x-1+u}\bmod p$ implies that $x$ is a fixed point on the main diagonal of the corresponding Costas array. Consequently, the number of solutions is equal to the number of fixed points on the main diagonal. Since the fixed points on the main diagonal form a Golomb ruler, via Lemma 1, the number of the fixed points is strictly less than
@@
Case (i): Suppose that $f$ and $g$ are both exponential Costas arrays. According to [DG2011Maximal, Theorem 1], we have
$\max\limits_{u}\max\limits_{\begin{subarray}{c}f,g\in\mathcal{W}_{p}\end{subarray}}\Psi_{f,g}(u,0)\leq 2.$
if $p$ is a safe prime, and
$\max\limits_{u}\max\limits_{\begin{subarray}{c}f,g\in\mathcal{W}_{p}\end{subarray}}\Psi_{f,g}(u,0)=\frac{p-1}{t}.$
otherwise.
Case (ii): Suppose that $f$ and $g$ are both power permutations. From [Ardalani2023Contributions, Theorem 5.7 and Theorem 5.9], it follows that, for $(u,v)=(0,0)$,
$\max\limits_{\begin{subarray}{c}f,g\in\mathcal{P}_{p}\\ f\neq g\end{subarray}}\Psi_{f,g}(0,0)=\frac{p-1}{t},$
@@
$\max\limits_{u\neq 0}\max\limits_{\begin{subarray}{c}f,g\in\mathcal{P}_{p}\\ f\neq g\end{subarray}}\Psi_{f,g}(u,0)\leq\left\lceil\frac{p-2}{p-1}(1+\sqrt{p})\right\rceil.$
@@
$\max\limits_{u}\max\limits_{\begin{subarray}{c}f,g\in\mathcal{P}_{p}\\ f\neq g\end{subarray}}\Psi_{f,g}(u,0)\leq\frac{p-2}{p-1}(1+\sqrt{p}),$
if $p$ is a safe prime, and
$\max\limits_{u}\max\limits_{\begin{subarray}{c}f,g\in\mathcal{P}_{p}\\ f\neq g\end{subarray}}\Psi_{f,g}(u,0)=\frac{p-1}{t},$
otherwise, since
$\frac{p-1}{t}\geq\sqrt{2(p-1)}\geq\frac{p-2}{p-1}(1+\sqrt{p})\text{ for }p\geq 11.$
@@
$\max\limits_{u}\max\limits_{\begin{subarray}{c}f,g\in\mathcal{PW}_{p}\end{subarray}}\Psi_{f,g}(u,0)\leq(p-1)^{1/2}+(p-1)^{1/4}+\frac{1}{2},$
if $p$ is a safe prime, and
@@
including the numerical results in Table V.
@@
For a prime $p\geq 5$, the family $\mathcal{PW}^{l}_{p}$ of order p1p-1, is defined as
@@
where $\mathcal{W}^{l}_{p}$ and $\mathcal{P}_{p}$ are defined in Construction 4 and Construction 5, respectively. The size of the family $\mathcal{PW}^{l}_{p}$ equals $2\phi(p-1)-1$, where $\phi$ is Euler’s totient function.